OpenStudy (anonymous):

Hey! I was hoping for someone to check my answer? I need to fill in the missing blanks to these two problems: 1. _ _ _(pi/3)sin(pi/6)=1/2(sin(pi/2)-sin(pi/6)) For the missing part, I thought the answer would just be sin?? 2. sin(pi/4)sin(pi/6)=1/2(_ _ _(pi/12)-cos(5pi/12)) I was thinking this one's missing part is: cos?? Help please!

OpenStudy (anonymous):

@ganeshie8 Help por favor? :)

ganeshie8 (ganeshie8):

You're using these formulas : http://prntscr.com/3rd6ss right ?

OpenStudy (anonymous):

Attempting too, yes... For some reason. Filling in the blanks is really confusing for me.. Too check myself.. I've been using this calculator.. Which probably isn't very helpful in learning how to solve... http://symbolab.com/solver/trigonometric-identity-calculator

ganeshie8 (ganeshie8):

I see... it will become easy if we break it down into 1-2 steps : _ _ _(pi/3)sin(pi/6)=1/2(sin(pi/2)-sin(pi/6)) is same as : _ _ _(pi/3)sin(pi/6)=1/2(sin(pi/3 + pi/6)-sin(pi/3 - pi/6))

OpenStudy (anonymous):

Oh! So I would solve the equation on the right to get the missing part on the left?

ganeshie8 (ganeshie8):

yes : sin(A+B) = sinAcosB + cosAsinB sin(A-B) = sinAcosB - cosAsinB ----------------------------------- sin(A+B) - sin(A-B) = 2cosAsinB

ganeshie8 (ganeshie8):

so you should fill the first blank by `cos`

OpenStudy (anonymous):

Oh!! That makes sense. so it'd look like: cos(pi/3)sin(pi/6)=1/2(sin(pi/2)-sin(pi/6))?

ganeshie8 (ganeshie8):

Yep !

OpenStudy (anonymous):

Yay! for cos being for the first!

ganeshie8 (ganeshie8):

second one is correct ! it has to be `cos`

ganeshie8 (ganeshie8):

http://prntscr.com/3rd6ss look up second identity ^

OpenStudy (anonymous):

Whoops Idk why I sent that again

OpenStudy (anonymous):

I was gonna ask.. Could you help me with using squared identities?

ganeshie8 (ganeshie8):

to be honest - i always get confuse with these product formulas pretty badly, and always derive them from the sum formulas when needed. We only need to remember four formulas : ``` sin(A+B) = sinAcosB + cosAsinB sin(A-B) = sinAcosB - cosAsinB cos(A+B) = cosAcosB - sinAsinB cos(A-B) = cosAcosB + sinAsinB ```

ganeshie8 (ganeshie8):

yeah sure, ask...

OpenStudy (anonymous):

Thank you for those formula's.. I saved them.. Okay so I'm supposed to use square identities to simplify sin^2sin^2x... I'm pretty lost.

OpenStudy (anonymous):

I have answer choices if that'd help a bit.

ganeshie8 (ganeshie8):

Is the expression just that ? \(\large \sin^2 (\sin^2x)\) ? yeah provide the answer choices...

OpenStudy (anonymous):

No, it just looks like: sin^2x sin^2x So it's just repeated.. It'll take me second to type them out! Don't run away!

ganeshie8 (ganeshie8):

I'll be here - may be take a snap and attach the file if possible

OpenStudy (anonymous):

K! Here it is!

ganeshie8 (ganeshie8):

we use below two formulas : \(\large \sin^2 x = \dfrac{1-\cos (2x)}{2}\) \(\large \cos^2 x = \dfrac{1+\cos (2x)}{2}\)

ganeshie8 (ganeshie8):

\(\large \sin^2x\times \sin^2x = \left(\sin^2x\right)^2 = \left(\dfrac{1-\cos(2x)}{2}\right)^2 \)

ganeshie8 (ganeshie8):

\(\large = \dfrac{1}{4}\left(1-\cos(2x)\right)^2 \)

ganeshie8 (ganeshie8):

use (a-b)^2 formula and expand...

ganeshie8 (ganeshie8):

\(\large = \dfrac{1}{4}\left(1-2\cos(2x) + \cos^2(2x)\right) \)

ganeshie8 (ganeshie8):

use the second formula for cos^2 term

ganeshie8 (ganeshie8):

\(\large = \dfrac{1}{4}\left(1-2\cos(2x) + \dfrac{1+\cos(4x)}{2}\right) \)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

Oh!! So I simplify that and it'll give me one of the answer choices? Seems simple enough! mmkay I'm gonna attempt to simplify. Just a sec

ganeshie8 (ganeshie8):

just pull out the 1/2 : \(\large = \dfrac{1}{8}\left(2-4\cos(2x) + 1+\cos(4x) \right) \)

OpenStudy (anonymous):

Oh, gotcha. So would that be sin^2x sin^2x simplified??? The reason I ask is because the answers I have to choose from are different.

ganeshie8 (ganeshie8):

add 2+1 ^

ganeshie8 (ganeshie8):

\(\large = \dfrac{1}{8}\left(3-4\cos(2x) +\cos(4x) \right) \)

OpenStudy (anonymous):

Ahh, so the process keeps going? Geez no wonder I am terrible at math. So, the answer would be B! (3-4cos(2x)+cos(4x))/8 !!!

ganeshie8 (ganeshie8):

lol you're very good at math.. yes B is the answer - lets verify it with wolfram : http://www.wolframalpha.com/input/?i=simplify+sin%5E2x*sin%5E2x

OpenStudy (anonymous):

Yay for wolfram!! You're so helpful, thank you for putting up with my slowness!

ganeshie8 (ganeshie8):

np, you're welcome :)